# The Jordan canonical form of the Fréchet derivative of a matrix function

### v2

@@ 1 Motivation and statement of the problem
Fix a pair of square matrices $X\in\mathbb{F}^{m\times m},Y\in\mathbb{F}^{n\times n}$. The set of all matrices of the form (3), equipped with matrix addition and matrix multiplication, is a commutative ring $\mathcal{K}_{X,Y}\subsetneq\mathbb{F}^{mn\times mn}$. Moreover, if we denote by $\mu_{M}(t)$ the minimal polynomial of the square matrix $M$, then by the first isomorphism theorem
@@ 2 Simplifying preliminaries
The problem is now to obtain the JCF of the bTTb matrix (5), or equivalently, as the only eigenvalue is obviously $p(\lambda,\mu)$, the ranks of the matrices $(P(J_{m}(\lambda),J_{n}(\mu))-p(\lambda,\mu)I_{mn})^{k}$, $k=1,\dots,\nu$, where $\nu$ the first value of $k$ for which the rank equals to 0. Since the value of $p(\lambda,\mu)$ is irrelevant to the answer, we may also assume with no loss of generality that it is 0, or in other words that $(P(J_{m}(\lambda),J_{n}(\mu))$ is nilpotent.
@@
Then $W$ and $Z_{m-1}$ are similar over $R$, i.e., they are similar and the similarity matrix can be taken having block elements in $R$.
@@
There is no loss of generality in assuming $A_{0}=0$. Indeed, we may equivalently solve $(Z_{m-1}-I_{m}\otimes A_{0})X=X(Z_{1}-I_{m}\otimes A_{0})$ because every matrix in $R^{m\times m}$ commutes with $I_{m}\otimes A_{0}$. Let $U=N_{m}\otimes I_{n}\in\hat{R}$ and, for $k=1,\dots,m-1$, define $Z_{k}=\sum_{i=1}^{k}A_{i}U^{i}$. (Note that this definition also covers $Z_{1},Z_{m-1}$.) The action of $U$ when acting from the left is an upwards row shift, while the action of $U$ from the right is a rightwards column shift. Since $R$ is a commutative ring we have [^1] Here and below the notation $X(a:b,c:d)$, where $a\leq b,c\leq d$ are positive integers, indicates the submatrix of $X$ obtained by selecting the rows with indices $a,\dots,b$ and the columns with indices $c,\dots,d$. $XZ_{1}=A_{1}\begin{bmatrix}0&X(1:m-1,1:m-1)\\ 0&0\end{bmatrix}$. On the other hand, $Z_{m-1}X=\sum_{i=1}^{m-1}A_{i}\begin{bmatrix}0&X(i+1:m,i+1:m)\\ 0&0\end{bmatrix}$ ( the blocks have different sizes for different terms in the summation). Therefore, $D_{ij}=0$ if ji<2j-i<2, while $D_{ij}=A_{j-i}-A_{1}X_{i,j-1}+\sum_{\ell=1}^{j-i-1}(A_{\ell}X_{i+\ell,j})$ if $j-i\geq 2$. Thus, $D_{ij}$ is an affine function $A_{1}$: $D_{ij}=D_{ij}^{(0)}+D_{ij}^{(1)}A_{1}$. In particular, when $j\geq i+2$, $D_{ij}^{(1)}=X_{i+1,j}-X_{i,j-1}$ and $D_{i,j}^{(0)}=\sum_{\ell=2}^{j-i-1}(A_{\ell}X_{i+\ell,j})$. Both $X_{i+1,j}$ and $X_{i,j-1}$ belong to the (ji1)(j-i-1)th superdiagonal [^2] The first superdiagonal of a square matrix is the set of the elements directly above the diagonal. Recursively, the $k$th superdiagonal is defined as the set of the elements directly above the (k1)(k-1)th superdiagonal. of $X$. Moreover, the unknown elements of $X$ appearing in $D^{(0)}_{ij}$ must belong to one of the first ji2j-i-2 superdiagonals of $X$. We conclude that the following procedure can be applied to solve the problem:
@@ 3.1 Generic case: Full solution
For each Jordan block of $X$ of size $m$ with eigenvalue $\lambda$ and for each Jordan block of $Y$ of size $n$ with eigenvalue $\mu$, assume that $(\lambda,\mu)$ is such that (6) holds. Then in the JCF of the matrix (3), there are, associated with each eigenvalue $p(\lambda,\mu)$:
@@
In fact, $p(x,y)=x+y$ as in Example 3.3 corresponds more generally to an important special case of Theorem 3.1: The JCF of a Kronecker sum. When specialized to this case, from Theorem 3.1 we recover a classical result that we recall below.
@@
Let $X=J_{2}(0)\oplus J_{1}(1)$ and $Y=J_{2}(2)\oplus J_{1}(3)$, where $\oplus$ denotes, here and throughout the paper, the direct sum of two matrices. We study the JCF of $P(X,Y)$ with $p(x,y)=y-2x+xy-y^{2}$. From the interaction the Jordan block with eigenvalue $2$ in $Y$ and any Jordan block in $X$, we get $p_{x}=0\neq p_{y}$; in the other cases we get $p_{x}\neq 0\neq p_{y}$. Carefully applying the appropriate cases in Theorem 3.1 to the four possible interactions of a Jordan block in $X$ with one in $Y$, we obtain that the JCF of $P(X,Y)=I_{3}\otimes Y-2X\otimes I_{3}+X\otimes Y-I_{3}\otimes Y^{2}$ is
@@ 3.1.1 The Jordan canonical form of $P(J_{m}(\lambda),J_{n}(\mu))$ for $p_{x}\neq 0$, $p_{y}\neq 0$
As discussed, this case appeared as [12, Corollary 1]. To make the paper self-contained, we nevertheless give an independent proof. Since $p_{x}\neq 0$, we can apply Lemma 2.4 and claim that it suffices to find the JCF of $I_{m}\otimes B+N_{m}\otimes I_{n}$, where $B$ is defined in (5). Then, since $p_{y}\neq 0$, we can apply Lemma 2.4 again to reduce $B$ to bidiagonal form via a certain similarity matrix $X$. Hence, by applying the similarity transformation associated with $I_{m}\otimes X$ to our larger matrix, we reduce to the problem of finding the JCF of
@@
The matrix above is a Kronecker sum, and it is known that the lengths of its Jordan chains are m+n+12km+n+1-2k for $k=1,2,\dots,\min(m,n)$. This is stated in [13, Theorem 2] (a crucial claim in [13] is verified in [11]). An alternative approach is discussed in [4], where it is shown that this problem is related to the study of the $k$-paths of the Cartesian product of $\Gamma_{m}$ and $\Gamma_{n}$, where $\Gamma_{k}$ is the digraph having vertices $1,2,\dots,k$ and arcs $12,23,\dots,(k-1)k$. We refer the reader to [4] for definitions and details. A slight modification to the proof [4, Lemma 4.3] leads to the same conclusion as in [13, Theorem 2]. We also note that, substituting every occurrence of $I_{n}$ in the first block superdiagonal with $I_{n}+N_{n}$ via a block-diagonal similarity, we could equivalently have used the results concerning Kronecker products: see [4, Lemma 4.3], and also [1, 9, 11, 13]. Therefore the first item in Theorem 3.1 is proved.
@@ 3.1.2 The Jordan canonical form of $P(J_{m}(\lambda),J_{n}(\mu))$ for $p_{x}\neq 0=p_{y}$
We now address the case where only one of the two first order derivatives is zero. There is no loss of generality in assuming that the nonzero one is $p_{x}$, as otherwise we can just swap the roles of the triples $(x,\lambda,m)$ and $(y,\mu,n)$ via Lemma 2.2. Following the same approach as above, we therefore aim to find the JCF of
@@
where $B$ is now an upper triangular Toeplitz matrix whose first superdiagonal is 0. We need a variation on the theme of Lemma 2.4.
@@
Let $\{A_{i}\}_{i=0}^{m-1}\subset R$. Define $Z_{m-1}\in\hat{R}$ as follows:
@@
Proof. We give a sketch of the proof, that follows a constructive strategy similar to the one of Lemma 2.4. One can find a similarity matrix $X\in R^{m\times m}$, upper triangular with all diagonal elements equal to $I_{n}$ and the first $r$ (block) rows equal to the rows of an identity matrix, such that $Z_{m-1}X=X(I_{m}\otimes A_{0}+N_{m}^{r}\otimes A_{r})$. It is then easy to find a (block) diagonal similarity matrix $D$ such that $D^{-1}(I_{m}\otimes A_{0}+N_{m}^{r}\otimes A_{r})D=(I_{m}\otimes A_{0}+N_{m}^{r}\otimes I_{n})$.
Using Lemma 3.6 (in the scalar case $R=\mathbb{F}$) we see that it suffices to compute the JCF of the nilpotent matrix $N=I_{m}\otimes N_{n}^{r}+N_{m}\otimes I_{n}$. Let $q(x,y)=x+y$; then $N=Q(N_{m},N_{n}^{r})$. Note that $q_{x}=q_{y}=1$ and therefore we can apply the results obtained in the previous section. The JCF of $N_{m}$ is $N_{m}$ itself, while the one of $N_{n}^{r}$ is provided by the following corollary. Corollary 3.7 is a straightforward consequence of [6, Theorem 1.36], or can also be verified by direct inspection:
@@ 3.2 Degenerate case: General bounds
The previous observation is frightening, and we will not delve further into the epic undertaking of a complete solution to the bivariate Jordan problem. Instead, we give some general bounds on the size and number of the Jordan blocks in the degenerate case when (6) fails, based on the the concept of local degree, that we define next. Given a bivariate polynomial $p(x,y)$ as in (2), we define the local degree of $p$ at $(\lambda,\mu)\in\mathbb{F}^{2}$ as the smallest positive integer $d\geq 1$ such that there exists a biindex $\alpha=(\beta,\gamma)\in\mathbb{N}^{2}$ with $|\alpha|=\beta+\gamma=d$ and $\hat{\partial}_{z^{\alpha}}p(\lambda,\mu)=\hat{\partial}_{x^{\beta}y^{\gamma}}p(\lambda,\mu)\neq 0$.
Recall Remark 3.2: If either $m=1$ or $n=1$ (or both), Theorem 3.1 can be applied regardless of the value of $p_{x}$ (if $m=1$) or regardless of the value of $p_{y}$ (if $n=1$). For this reason, throughout the present section we assume $m,n>1$.
@@
Let $d\geq 1$ be the local degree of $p(x,y)$ at $(\lambda,\mu)$, and write $P(J_{m}(\lambda),J_{n}(\mu))=p(\lambda,\mu)I_{nm}+Z_{m-1}$, where $Z_{m-1}$ is nilpotent. Then the size of the Jordan blocks of $P(J_{m}(\lambda),J_{n}(\mu))$ is bounded above by $\lceil\frac{m+n-1}{d}\rceil$.
Proof. We will show that $Z_{m-1}^{k}=0$ for $k\geq\lceil\frac{m+n-1}{d}\rceil$. Recall from Section 2 that the commutative ring of $m\times m$ triangular Toeplitz matrices with elements in a commutative ring $R$ is isomorphic to the quotient ring $R[y]/\langle y^{m}\rangle$. Therefore, in order to analyze a power of $Z_{m-1}$ it suffices to consider the scalar polynomial $(A_{0}+A_{1}y+\dots+A_{m-1}y^{m-1})^{k}\bmod y^{m}$. Before quotienting, by the multinomial theorem we have
$(A_{0}+A_{1}y+\dots+A_{m-1}y^{m-1})^{k}=\sum_{|\alpha|=k}\frac{k!}{\alpha_{0}!\alpha_{1}!\cdots\alpha_{m-1}!}\prod_{i=0}^{m-1}A_{i}^{\alpha_{i}}y^{i\alpha_{i}},$
where $\alpha=(\alpha_{0},\dots,\alpha_{m-1})$ is a multiindex equipped with the usual multiindex $1$-norm. Let us study the term in $y^{h}$ in the expansion above, $h\leq m-1$. By assumption, $\prod_{j=0}^{d-1}A_{j}^{\alpha_{j}}=0$ as long as $\sum_{j=0}^{d-1}(d-j)\alpha_{j}\geq n$. We now aim to minimize the value of $\sum_{j=0}^{d-1}(d-j)\alpha_{j}$ subject to $|\alpha|=k$ and $\sum_{i=0}^{m-1}i\alpha_{i}=h$. By the constraints we obtain
@@
Hence, when $dk\geq m+n-1$, all the coefficients of $y^{h}$ are zero for all $h=0,\dots,m-1$.
@@
When applied to Jordan blocks, [10, Corollary 4.3] is a special case (for $d=1$) of Proposition 3.8.
@@
Proof. Write $P(J_{m}(\lambda),J_{n}(\mu))=p(\lambda,\mu)I_{mn}+Z_{m-1}$. If $d\geq m+n-1$, then every nonzero addend in $Z_{m-1}$ is a multiple of $N_{m}^{a}\otimes N_{n}^{b}$ with $a+b\geq m+n-1$. If $a\geq m$, $N_{m}^{a}=0$; and if $a<m$ then b>n1b>n-1 and $N_{n}^{b}=0$. Hence, $Z_{m-1}=0$.
Suppose now d<n+m1d<n+m-1. Recalling Lemma 2.2, we may also assume with no loss of generality $m\leq n$. Since the number of Jordan blocks is the nullity of $Z_{m-1}$, our goal is to prove $\mathfrak{b}\leq\dim\ker Z_{m-1}\leq\mathfrak{a}$.
@@
  • Upper bound. If $d\geq\min(m,n)$, then $\mathfrak{a}=mn$ and thus the statement is trivial; hence, we assume $d<\min(m,n)$. Since the number of Jordan blocks is the nullity of $Z_{m-1}$, in that case (recalling also $m\leq n$) we aim to prove $\mathrm{rank}(Z_{m-1})\geq n(m-d)$.
    By Lemma 3.10, for all $k$, $Z_{m-1}$ maps $V_{k}/V_{k-1}$ to $V_{k-d}$, where if $k<d$ we define $V_{k-d}=V_{0}=\{0\}$. Let us permute rows and columns of $Z_{m-1}$, reordering them by increasing $V$-degree (for equal $V$-degrees, we keep the original order; in other words, after the permutation we order rows and columns by the grlex monomial order: $e_{i_{1}}\otimes f_{j_{1}}<e_{i_{2}}\otimes f_{j_{1}}$ when either $i_{1}+j_{1}<i_{2}+j_{2}$ or $i_{1}+j_{1}=i_{2}+j_{2}$ and $i_{1}<i_{2}$). We thus obtain a new matrix $T$ which is block upper triangular. Furthermore, both the block diagonal and the first d1d-1 block superdiagonals of $T$ are zero. Finally, $\mathrm{rank}(Z_{m-1})=\mathrm{rank}(T)\geq\sum_{k=d+1}^{m+n-1}\mathrm{rank}(R_{k})$ where $R_{k}$ are the blocks on the $(d+1)$-th block superdiagonal of $T$. In other words, $R_{k}\in\mathbb{F}^{u_{k-d}\times u_{k}}$ is the banded (with band size $d+1$) Toeplitz rectangular matrix that represents the linear map $V_{k}/V_{k-1}\rightarrow V_{k-d}/V_{k-d-1}$, obtained by appropriately restricting the action of $Z_{m-1}$. We know that at least one diagonal in the Toeplitz matrices $R_{k}$ is nonzero, and the ranks of $R_{k}$ are minimal when there is only one nonzero diagonal, and it is in an extremal position. In particular, since $m\leq n$, the worst-case scenario occurs when $\hat{\partial}_{x^{d}}p(\lambda,\mu)\neq 0$ but the other derivatives of order $d$ are zero. Since we eventually have to sum these lower bounds over $k$,

### v3

@@ 1 Motivation and statement of the problem
Fix a pair of square matrices $X\in\mathbb{F}^{m\times m},Y\in\mathbb{F}^{n\times n}$. The set of all matrices of the form (3), equipped with matrix addition and matrix multiplication, is a commutative ring $\mathcal{K}_{X,Y}\subseteq\mathbb{F}^{mn\times mn}$. Moreover, if we denote by $\mu_{M}(t)$ the minimal polynomial of the square matrix $M$, then by the first isomorphism theorem
@@ 2 Simplifying preliminaries
The problem is now to obtain the JCF of the bTTb matrix (5), or equivalently, as the only eigenvalue is obviously $p(\lambda,\mu)$, the ranks of the matrices $(P(J_{m}(\lambda),J_{n}(\mu))-p(\lambda,\mu)I_{mn})^{k}$, $k=1,\dots,\nu$, where $\nu$ the smallest value of $k$ for which the rank equals to 0. Since the value of $p(\lambda,\mu)$ is irrelevant to the answer, we may also assume with no loss of generality that it is 0, or in other words that $(P(J_{m}(\lambda),J_{n}(\mu))$ is nilpotent.
@@
Then $W$ and $Z_{m-1}$ are similar over $R$, i.e., they are similar and the similarity matrix can be taken to have block elements in $R$.
@@
There is no loss of generality in assuming $A_{0}=0$. Indeed, we may equivalently solve $(Z_{m-1}-I_{m}\otimes A_{0})X=X(Z_{1}-I_{m}\otimes A_{0})$ because every matrix in $R^{m\times m}$ commutes with $I_{m}\otimes A_{0}$. Let $U=N_{m}\otimes I_{n}\in\hat{R}$ and, for $k=1,\dots,m-1$, define $Z_{k}=\sum_{i=1}^{k}A_{i}U^{i}$. (Note that this definition also covers $Z_{1},Z_{m-1}$.) The action of $U$ when acting from the left is an upwards row shift, while the action of $U$ from the right is a rightwards column shift. Since $R$ is a commutative ring we have [^1] Here and below the notation $X(a:b,c:d)$, where $a\leq b,c\leq d$ are positive integers, indicates the submatrix of $X$ obtained by selecting the rows with indices $a,\dots,b$ and the columns with indices $c,\dots,d$. $XZ_{1}=A_{1}\begin{bmatrix}0&X(1:m-1,1:m-1)\\ 0&0\end{bmatrix}$. On the other hand, $Z_{m-1}X=\sum_{i=1}^{m-1}A_{i}\begin{bmatrix}0&X(i+1:m,i+1:m)\\ 0&0\end{bmatrix}$ ( the blocks have different sizes for different terms in the summation). Therefore, $D_{ij}=0$ if ji<2j-i<2, while $D_{ij}=A_{j-i}-A_{1}X_{i,j-1}+\sum_{\ell=1}^{j-i-1}(A_{\ell}X_{i+\ell,j})$ if $j-i\geq 2$. Thus, $D_{ij}$ is an affine function of $A_{1}$: $D_{ij}=D_{ij}^{(0)}+D_{ij}^{(1)}A_{1}$. In particular, when $j\geq i+2$, $D_{ij}^{(1)}=X_{i+1,j}-X_{i,j-1}$ and $D_{i,j}^{(0)}=\sum_{\ell=2}^{j-i-1}(A_{\ell}X_{i+\ell,j})$. Both $X_{i+1,j}$ and $X_{i,j-1}$ belong to the (ji1)(j-i-1)th superdiagonal [^2] The first superdiagonal of a square matrix is the set of the elements directly above the diagonal. Recursively, the $k$th superdiagonal is defined as the set of the elements directly above the (k1)(k-1)th superdiagonal. of $X$. Moreover, the unknown elements of $X$ appearing in $D^{(0)}_{ij}$ must belong to one of the first ji2j-i-2 superdiagonals of $X$. We conclude that the following procedure can be applied to solve the problem:
@@ 3.1 Generic case: Full solution
For each Jordan block of $X$ of size $m$ with eigenvalue $\lambda$ and for each Jordan block of $Y$ of size $n$ with eigenvalue $\mu$, assume that $(\lambda,\mu)$ is such that (6) holds. Then in the JCF of the matrix (3), there are, associated with the eigenvalue $p(\lambda,\mu)$:
@@
In fact, $p(x,y)=x+y$ as in Example 3.3 corresponds more generally to an important special case of Theorem 3.1: The JCF of a Kronecker sum. When specializing to this case, from Theorem 3.1 we recover a classical result that we recall below.
@@
Let $X=J_{2}(0)\oplus J_{1}(1)$ and $Y=J_{2}(2)\oplus J_{1}(3)$, where $\oplus$ denotes, here and throughout the paper, the direct sum of two matrices. We study the JCF of $P(X,Y)$ with $p(x,y)=y-2x+xy-y^{2}$. From the interaction of the Jordan block with eigenvalue $2$ in $Y$ and any Jordan block in $X$, we get $p_{x}=0\neq p_{y}$; in the other cases we get $p_{x}\neq 0\neq p_{y}$. Carefully applying the appropriate cases in Theorem 3.1 to the four possible interactions of a Jordan block in $X$ with one in $Y$, we obtain that the JCF of $P(X,Y)=I_{3}\otimes Y-2X\otimes I_{3}+X\otimes Y-I_{3}\otimes Y^{2}$ is
@@ 3.1.1 The Jordan canonical form of $P(J_{m}(\lambda),J_{n}(\mu))$ for $p_{x}\neq 0$, $p_{y}\neq 0$
As discussed, this case appeared as [12, Corollary 1]. To make the paper self-contained, we nevertheless give an independent proof. Since $p_{x}\neq 0$, we can apply Lemma 2.4 and claim that it suffices to find the JCF of $I_{m}\otimes B+N_{m}\otimes I_{n}$, where $B$ is defined in (5). Then, since $p_{y}\neq 0$, we can apply Lemma 2.4 again to reduce $B$ to bidiagonal form via a certain similarity matrix $X$. Hence, we reduce to the problem of finding the JCF of
@@
The matrix above is a Kronecker sum, and it is known that the lengths of its Jordan chains are m+n+12km+n+1-2k for $k=1,2,\dots,\min(m,n)$. This is stated in [13, Theorem 2] (a crucial claim in [13] is verified in [11]). An alternative approach is discussed in [4], where it is shown that this problem is related to the study of the $k$-paths of the Cartesian product of $\Gamma_{m}$ and $\Gamma_{n}$, where $\Gamma_{k}$ is the digraph having vertices $1,2,\dots,k$ and arcs $12,23,\dots,(k-1)k$. We refer the reader to [4] for definitions and details. A slight modification to the proof of [4, Lemma 4.3] leads to the same conclusion as in [13, Theorem 2]. We also note that, substituting every occurrence of $I_{n}$ in the first block superdiagonal with $I_{n}+N_{n}$ via a block-diagonal similarity, we could equivalently have used the results concerning Kronecker products: see [4, Lemma 4.3], and also [1, 9, 11, 13]. Therefore the first item in Theorem 3.1 is proved.
@@ 3.1.2 The Jordan canonical form of $P(J_{m}(\lambda),J_{n}(\mu))$ for $p_{x}\neq 0=p_{y}$
We now address the case where only one of the two first order Hasse derivatives is zero. There is no loss of generality in assuming that the nonzero one is $p_{x}$, as otherwise we can just swap the roles of the triples $(x,\lambda,m)$ and $(y,\mu,n)$ via Lemma 2.2. Following the same approach as above, we therefore aim to find the JCF of
@@
where $B$ is now an upper triangular Toeplitz matrix whose first superdiagonal is 0. We now need a variation on the theme of Lemma 2.4.
@@
Let $\{A_{i}\}_{i=0}^{m-1}\subset R$. Define $Z_{m-1}\in\hat{R}$ by prescribing its first row to have elements $A_{0},0=A_{1}=\dots=A_{r-1},A_{r},\dots,A_{m-1}$. That is,
@@
Proof. We sketch the proof, that is constructive and similar to the one of Lemma 2.4. One can find a similarity matrix $X\in R^{m\times m}$, upper triangular with all diagonal elements equal to $I_{n}$ and the first $r$ (block) rows equal to the rows of an identity matrix, such that $Z_{m-1}X=X(I_{m}\otimes A_{0}+N_{m}^{r}\otimes A_{r})$. It is then easy to find a (block) diagonal similarity matrix $D$ such that $D^{-1}(I_{m}\otimes A_{0}+N_{m}^{r}\otimes A_{r})D=(I_{m}\otimes A_{0}+N_{m}^{r}\otimes I_{n})$.
Using Lemma 3.6 (in the scalar case $R=\mathbb{F}$) we see that it suffices to compute the JCF of the nilpotent matrix $N=I_{m}\otimes N_{n}^{r}+N_{m}\otimes I_{n}$. Let $q(x,y)=x+y$; then $N=Q(N_{m},N_{n}^{r})$. Note that $q_{x}=q_{y}=1$ and therefore we can apply the results obtained in the previous section. The JCF of $N_{m}$ is $N_{m}$ itself, while the one of $N_{n}^{r}$ is provided by Corollary 3.7, which is a straightforward consequence of [6, Theorem 1.36].
@@ 3.2 Degenerate case: General bounds
The previous observation is frightening, and we will not delve further into the epic undertaking of a complete solution to the bivariate Jordan problem. Instead, we give some general bounds on the size and number of the Jordan blocks in the degenerate case when (6) fails, based on the the concept of local degree, that we define next. Given a bivariate polynomial $p(x,y)$ as in (2), we define the local degree of $p$ at $(\lambda,\mu)\in\mathbb{F}^{2}$ as the smallest positive integer $d\geq 1$ such that there exists a biindex $\alpha=(\beta,\gamma)\in\mathbb{N}^{2}$ with $|\alpha|=\beta+\gamma=d$ and $\hat{\partial}_{z^{\alpha}}p(\lambda,\mu)=\hat{\partial}_{x^{\beta}y^{\gamma}}p(\lambda,\mu)\neq 0$. Note that $d\in\mathbb{N}$ except if all the Hasse derivatives of $p$ are zero, or equivalently, except if $p$ is constant. As the bivariate Jordan problem is trivial when $p$ is a constant, we henceforth exclude this case.
Recall Remark 3.2: If either $m=1$ or $n=1$ (or both), Theorem 3.1 can be applied regardless of the value of $p_{x}$ (if $m=1$) or regardless of the value of $p_{y}$ (if $n=1$). For this reason, from now on we also assume $m,n>1$.
@@
Suppse that $p(x,y)$ is not constant and let $d\geq 1$ be the local degree of $p(x,y)$ at $(\lambda,\mu)$. Write $P(J_{m}(\lambda),J_{n}(\mu))=p(\lambda,\mu)I_{nm}+Z_{m-1}$, where $Z_{m-1}$ is nilpotent. Then the size of the Jordan blocks of $P(J_{m}(\lambda),J_{n}(\mu))$ is bounded above by $\lceil\frac{m+n-1}{d}\rceil$.
Proof. We will show that $Z_{m-1}^{k}=0$ for $k\geq\lceil\frac{m+n-1}{d}\rceil$. Recall from Section 2 that the commutative ring of $m\times m$ triangular Toeplitz matrices with elements in a commutative ring $R$ is isomorphic to the quotient ring $R[x]/\langle x^{m}\rangle$. Therefore, in order to analyze a power of $Z_{m-1}$ it suffices to consider the scalar polynomial $(A_{0}+A_{1}x+\dots+A_{m-1}x^{m-1})^{k}\bmod x^{m}$. Before quotienting, by the multinomial theorem we have
$(A_{0}+A_{1}y+\dots+A_{m-1}x^{m-1})^{k}=\sum_{|\alpha|=k}\frac{k!}{\alpha_{0}!\alpha_{1}!\cdots\alpha_{m-1}!}\prod_{i=0}^{m-1}A_{i}^{\alpha_{i}}x^{i\alpha_{i}},$
where $\alpha=(\alpha_{0},\dots,\alpha_{m-1})$ is a multiindex equipped with the usual multiindex $1$-norm. Let us study the term in $x^{h}$ in the expansion above, $h\leq m-1$. By definition of $d$ and (5), and since $R\cong\mathbb{F}[y]/\langle y^{n}\rangle$, it holds for all $j=0,\dots,d-1$ that $A_{j}\in R$ corresponds to an element $p_{j}]\in\langle[y]^{d-j}\rangle\subset\mathbb{F}[y]/\langle y^{n}\rangle$. Therefore, $\prod_{j=0}^{d-1}A_{j}^{\alpha_{j}}=0$ as long as $\sum_{j=0}^{d-1}(d-j)\alpha_{j}\geq n$. We now aim to minimize the value of $\sum_{j=0}^{d-1}(d-j)\alpha_{j}$ subject to $|\alpha|=k$ and $\sum_{i=0}^{m-1}i\alpha_{i}=h$. By the constraints we obtain
@@
Hence, when $dk\geq m+n-1$, all the coefficients of $x^{h}$ are zero for all $h=0,\dots,m-1$.
@@
Taking $d=1$ in Proposition 3.8 yields [10, Corollary 4.3].
@@
Proof. Write $P(J_{m}(\lambda),J_{n}(\mu))=p(\lambda,\mu)I_{mn}+Z_{m-1}$. If $d\geq m+n-1$, then by (8) every nonzero addend in $Z_{m-1}$ is a multiple of $N_{m}^{a}\otimes N_{n}^{b}$ with $a+b\geq m+n-1$. If $a\geq m$, $N_{m}^{a}=0$; and if $a<m$ then b>n1b>n-1 and $N_{n}^{b}=0$. Hence, $Z_{m-1}=0$.
Suppose now d<n+m1d<n+m-1. Recalling Lemma 2.2, we may also assume with no loss of generality $m\leq n$. Since the number of Jordan blocks in the JCF of $P(J_{m}(\lambda),J_{n}(\mu))$ is the nullity of $Z_{m-1}$, our goal is to prove $\mathfrak{b}\leq\dim\ker Z_{m-1}\leq\mathfrak{a}$.
@@
  • Upper bound. If $d\geq\min(m,n)$, then $\mathfrak{a}=mn$ and thus the statement is trivial; hence, we assume $d<\min(m,n)$. Since the number of Jordan blocks is the nullity of $Z_{m-1}$, in that case (still assuming $m\leq n$) we aim to prove $\mathrm{rank}(Z_{m-1})\geq n(m-d)$.
    By Lemma 3.10, for all $k$, $Z_{m-1}$ maps $V_{k}/V_{k-1}$ to $V_{k-d}$, where if $k<d$ we define $V_{k-d}=V_{0}=\{0\}$. Let us permute rows and columns of $Z_{m-1}$, reordering them by increasing $V$-degree (for equal $V$-degrees, we keep the original order; in other words, after the permutation we order rows and columns by the grlex monomial order: $e_{i_{1}}\otimes f_{j_{1}}<e_{i_{2}}\otimes f_{j_{1}}$ when either $i_{1}+j_{1}<i_{2}+j_{2}$ or $i_{1}+j_{1}=i_{2}+j_{2}$ and $i_{1}<i_{2}$). We thus obtain a new matrix $T$ which is block upper triangular [^3] Note that, generally, the blocks in a given superdiagonal of $T$ do not all have the same size.. Furthermore, both the main block diagonal and the first d1d-1 block superdiagonals of $T$ are zero. Finally, $\mathrm{rank}(Z_{m-1})=\mathrm{rank}(T)\geq\sum_{k=d+1}^{m+n-1}\mathrm{rank}(R_{k})$ where $R_{k}$ are the blocks on the $(d+1)$-th block superdiagonal of $T$. In other words, $R_{k}\in\mathbb{F}^{u_{k-d}\times u_{k}}$ is the banded (with band size $d+1$) Toeplitz rectangular matrix that represents the linear map $V_{k}/V_{k-1}\rightarrow V_{k-d}/V_{k-d-1}$, obtained by appropriately restricting the action of $Z_{m-1}$. We know that at least one diagonal in the Toeplitz matrices $R_{k}$ is nonzero, and the ranks of $R_{k}$ are minimal when there is only one nonzero diagonal, and it is in an extremal position. In particular, since $m\leq n$, the worst-case scenario occurs when $\hat{\partial}_{x^{d}}p(\lambda,\mu)\neq 0$ but the other derivatives of order $d$ are zero. Since we eventually have to sum these lower bounds over $k$,